OpenStudy (anonymous):

Find the product and simplify if possible. 5x/7w^4.w/8x

OpenStudy (saifoo.khan):

\[\frac{5x}{7w^4} \times \frac{w}{8x}\]

OpenStudy (saifoo.khan):

Like that?

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

\[\frac{5}{7w^3} \times \frac 1 8\]

OpenStudy (saifoo.khan):

\[\frac{5}{56w^3}\]